OpenStudy (xapproachesinfinity):

how would you show this series converge not using limit comparison test \[\sum_{k=1}^{\infty}\frac{3\sqrt{k}+2}{2k^2+5}\]

OpenStudy (xapproachesinfinity):

we could compare it with \[\frac{1}{k^{3/2}}\] but i want a systematic way rather than look for a p series that is greater and cvg

OpenStudy (amistre64):

well im thinking if we simplify it to k + 2 - 5k/sqrt(k) does this work simpler?

OpenStudy (amistre64):

5sqrt(k)/k^2 that is

OpenStudy (xapproachesinfinity):

hmm that would be good but i don't where did you get that lol

OpenStudy (amistre64):

multiply top and bottom by ksqrt(k) and then add zero to the top really

OpenStudy (xapproachesinfinity):

oh i see, i was just about thinking that is where you got it

OpenStudy (xapproachesinfinity):

the series that we get is \[\frac{5}{k^{3/4}}\]

OpenStudy (xapproachesinfinity):

oh yes 3/2, don't what is wrong with me today

OpenStudy (amistre64):

\[\frac{k\sqrt k}{k\sqrt k}\frac{3\sqrt{k}+2}{(2k^2+5)}\] \[\frac{3k^2+2k\sqrt k}{k\sqrt k(2k^2+5)}\] \[\frac{3k^2-(k^2+k^2)+2k\sqrt k+(5-5)}{k\sqrt k(2k^2+5)}\] \[\frac{2k^2+k^2+2k\sqrt k+5-5}{k\sqrt k(2k^2+5)}\] \[\frac{(2k^2+5)+(k^2+2k\sqrt k-5)}{k\sqrt k(2k^2+5)}\] hmm, seems i errored along the way

OpenStudy (xapproachesinfinity):

i actually did take \[\frac{3\sqrt{k}}{2k^2}=\frac{3}{2k^{3/2}}\] i just did some wrong math and made it power 3/4 which gave something unwanted

OpenStudy (amistre64):

:) bad math happens

OpenStudy (xapproachesinfinity):

darn it! too many bad math today!

OpenStudy (xapproachesinfinity):

it made think that the prof used something else lol

OpenStudy (xapproachesinfinity):

question does harmonic series appear to converge to zero? if we take out the first term a1=1

OpenStudy (xapproachesinfinity):

does not*

OpenStudy (amistre64):

harmonic does not converge no matter how many discrete terms you remove. infinity - 1 = infinity right?

OpenStudy (xapproachesinfinity):

yes true

OpenStudy (amistre64):

unless, you remove specific terms only get 1 + 1/2 + 1/4 + 1/8 + ... :) and the like

OpenStudy (xapproachesinfinity):

well does not \[1/2+1/3+1/4...................\] appear to be converging

OpenStudy (amistre64):

the sequence 1/n converges, but a convergent sequence does not necessarily gives us a convergent sum of the terms of a sequence

OpenStudy (xapproachesinfinity):

or may be I am perceiving this the wrong way

OpenStudy (xapproachesinfinity):

oh yeah i gotcha you i mixed up with sequence convergence :)

OpenStudy (amistre64):

if the list of terms (the sequence) does NOT go to zero, then we cannot converge the sum the converse of this is not true ..

OpenStudy (xapproachesinfinity):

i recall a youtube prof mentioned that it appear we will take forever to read the wall when he talked about this series

OpenStudy (xapproachesinfinity):

to reach the wall*

OpenStudy (amistre64):

1/2 = 1/2! 1/2 + 1/3 = 5/3! 5/6 + 1/4 = 21/4! 21/24 + 1/5 = 106/5! 106/5! + 1/6 = 636/6! dunno if theres a pattern that we could determine in the partial sums that we could then limit out to show its not convergent

OpenStudy (xapproachesinfinity):

hmm seems it is growing to infinity but slowly?

OpenStudy (amistre64):

yep